Write a function which is differentiable at infinite many points as well as non-differentiable at infinite many points.
1
Expert's answer
2021-10-12T14:54:27-0400

a function which is differentiable at infinite many points:

"f(x)=e^x" is differentiable at "x\in (\infin,\infin)"


a function which is non-differentiable at infinite many points:

"f(x)=tanx" has not derivative at "x=\pi /2+\pi n,\ n\isin Z"